# Factorized Motion Fields for Fast Sparse Input Dynamic View Synthesis

### v1

@@ 4.1. Evaluation Setup
Datasets: We evaluate our model on two popular multi-view dynamic scene datasets, namely N3DV (Pumarola et al., 2021) and InterDigital (Sabater et al., 2017) with two, three, and four input views. Following prior work (Fridovich-Keil et al., 2023), we downsample the videos spatially by a factor of two for all the experiments. We use the video at the center of the camera rig for testing and uniformly sample train videos from the remaining videos. N3DV dataset contains six real-world scenes with 17–21 static cameras per scene and 300 frames per viewpoint. The videos have a spatial resolution of $1352\times 1014$ and a frame rate of 30fps. InterDigital dataset contains multiple real-world scenes with 16 static cameras and varying number of frames per scene. We undistort the video frames using the radial distortion parameters provided with the dataset and use the undistorted videos for our experiments. We select five scenes that contain at least 300 frames and choose the first 300 frames. The videos have a spatial resolution of $1024\times 544$ and a frame rate of 30fps spanning 10 seconds in all our experiments.
@@
Table 3. Quantitative results on ablations: We evaluate the significance of the two flow priors we employ, namely the sparse flow prior ($\mathcal{P}_{\text{sf}}$) and the dense flow prior ($\mathcal{P}_{\text{df}}$) by disabling one component at a time. We also show the performance of the DyRF model without any priors for reference. We report the PSNR, SSIM, LPIPS scores for the rendered images and the depth MAE for the rendered depth maps on the N3DV dataset.
@@ 4.2. Comparisons and Implementation Details
We mainly compare the performance of our model against K-Planes. We use the official code released by the authors and modify it to implement our model. We refer to our base model without any priors as DyRF model. To test the superiority of our priors, we also compare our model against K-Planes with sparse depth priors and DyRF with dense flow priors across cameras. Since we use two factorized models instead of one used in K-Planes, we reduce the feature dimension in both of our models by half to keep the total number of parameters comparable to K-Planes. We train all the models for 30k iterations on a single NVIDIA RTX 2080 GPU. We randomly pick $s\in\{t-10,t+10\}$ to impose the flow prior losses and set $\lambda_{\text{sf}}=1$ and $\lambda_{\text{df}}=1$. We use the same values as suggested by K-Planes for all the remaining hyperparameters.
@@
Table 4. We test the performance of our DyRF model with dense flow priors across cameras and our reliable flow priors. The performance of the DyRF model without any priors is also shown for reference.
@@ 4.3. Results
We show the quantitative performance of our model and K-Planes in Tab. 2, where we observe that our ReFlow-DyRF model outperforms K-Planes across all the settings on both the datasets. We observe that the performance of all the models is relatively higher on the N3DV dataset as compared to the InterDigital dataset. This is perhaps due to the InterDigital dataset having larger motion and highly textured regions. From Figs. 6, 7 and 9, we observe that our model is able to correct errors both in moving objects and in static regions over K-Planes. In particular, we observe that K-Planes suffer from blur, deformation, and disappearance of moving objects while rendering novel views, which are mitigated by our model. The improvements are more starkly visible in the supplementary videos. We compare the rendered depth from both the models in Figs. 11 and 11 and observe that our model is able to reconstruct the depth more accurately than K-Planes. Tab. 2 also shows that our DyRF model without any priors performs better than K-Planes in the scenes with larger motion, such as the scenes in the InterDigital dataset, perhaps due to the temporal consistency enforced by the canonical volume. This is also visible in the example shown in Fig. 4. In our experiments, ReFlow-DyRF roughly took 1.5 hours and 5GB GPU memory to train on a single scene, and 6 seconds to render a single frame. The size of the saved model parameters is approximately 280MB. Please refer to Fridovich-Keil et al. (2023) for the training time for various models including K-Planes and DyNeRF.
Figure 5. Qualitative examples for depth priors vs flow priors: We observe that sparse depth prior is not very effective in improving the reconstruction quality when imposed on both K-Planes and our DyRF model. However, our sparse flow prior is highly effective in mitigating the distortions, while the use of sparse and dense flow priors in our final model leads to the best reconstruction quality.
Ablations We analyze the significance of each component of our model by disabling one component at a time in Tab. 3. We observe that the sparse flow prior is the most significant component of our model, giving a large boost in performance when included and a large drop in performance when excluded. The effect of sparse flow prior is more pronounced in terms of depth MAE, which measures the accuracy of the reconstructed 3D scene. This demonstrates the importance of reliable motion priors across cameras, even if sparse. Fig. 8 shows that sparse flow prior is more helpful in mitigating motion related artifacts such as motion blur, while within-camera dense flow prior mainly helps reduce distortions in the static regions. The combination of the two priors leads to the best performance by mitigating both artifacts.
@@
DyRF 0.81 0.39 0.68 0.22
DyRF + $\mathcal{P}_{\text{sd}}$ 0.80 0.34 0.70 0.22
DyRF + $\mathcal{P}_{\text{sf}}$ 0.87 0.20 0.73 0.13
ReFlow-DyRF 0.88 0.16 0.74 0.12
Table 5. We analyze the impact of our sparse flow prior that regularizes both the scene geometry and the learned motion against the sparse depth prior (Deng et al., 2022) that regularizes only the scene geometry. For reference, we also show the performance of our ReFlow-DyRF model. The models are trained with three input views on both the datasets.
Dense cross-camera flow priors vs our priors: To validate our hypothesis that a combination of cross-camera sparse flow and within-camera dense flow priors is more effective than dense flow priors across cameras, we impose cross-camera dense flow priors on our base DyRF model and evaluate the performance on both datasets. From Tab. 4 and Fig. 13, we observe that imposing noisy cross-camera dense flow priors leads to a large drop in performance, whereas our priors help improve the performance significantly.
Depth vs flow priors: We now analyze the effect of flow priors against that of depth priors in dynamic view synthesis with sparse input views. The depth priors guide the model on the exact locations of the objects in the scene, but do not account for the correlation of object locations across time instants. On the other hand, the flow priors account for such inter-frame correlations as well as encompass the information provided by the depth priors in static regions. To experimentally validate our hypothesis, we compare the performance of K-Planes and our DyRF models when regularized with sparse depth priors $\mathcal{P}_{\text{sd}}$ obtained through Colmap sparse reconstruction (Deng et al., 2022). We obtain $\mathcal{P}_{\text{sd}}$ for every frame and use them to constrain the K-Planes model as ${\mathcal{L}_{\text{sd}}}=\|z-\hat{z}\|^{2}$, where $z$ is the depth obtained in Eq. 6 and $\hat{z}$ is the depth prior. From Tab. 5, we observe that sparse flow priors lead to a significant improvement in performance as compared to that of sparse depth priors. The difference in performance can be more clearly observed in Fig. 5 and the supplementary videos. This shows the utility of motion priors in dynamic view synthesis with sparse input views.
Performance with increasing viewpoints: To understand the impact of our priors as the number of input views increases, we compare the performance of ReFlow-DyRF against K-Planes with different number of cameras on a single scene from N3DV dataset in Fig. 13. Our model significantly improves the performance over K-Planes in sparse viewpoint scenarios, while matching the performance of K-Planes with dense input views. Thus, we conclude that our priors are particularly helpful in improving the reconstruction of under-observed regions in the scene.
@@
DyRF 29.95 0.95 0.12 28.14 0.95 0.09
Table 6. We test the performance of our DyRF model against K-Planes with dense input views. We find that our approach of learning the motion model with a factorized volume is competitive with K-Planes. Please see the videos in the supplementary for visual comparisons.
Performance with dense input views: We now test the ability of our model in reconstructing the dynamic scene when provided with dense multi-view inputs. Since motion priors may not be needed when dense views are available, we analyze the performance of our base DyRF model without any priors. From Tab. 6, we observe that our DyRF model is competitive with K-Planes. Thus, one could employ our model in both dense and sparse input scenarios by additionally employing reliable flow priors in the latter case.
@@ 4.4. Limitations and Future Work
Our approach of mapping the scene at every time instant to the canonical volume implies that only the objects present in the canonical volume can be rendered. In other words, our model may not handle the case of objects entering or leaving the scene, which could happen over longer durations. This could be resolved by learning multiple models, each trained on self-contained shorter duration videos. We employ Colmap to generate sparse flow priors in our framework, which is a time-consuming process. In our experiments, we found that the generation of sparse flow priors takes about 45 minutes per scene. There is a need to explore faster approaches to determine reliable sparse correspondences across cameras. Further, our approach requires the cameras to be calibrated and time-synchronized apriori. It would be interesting to optimize the scene, camera parameters, and time-synchronization jointly in a single framework.
@@ 5. Conclusion
We consider the setting of fast dynamic view synthesis when only a few videos of the scene as observed from different static cameras are available. By exploiting the spatio-temporal correlation of the motion field, we design an explicit motion model using factorized representations that is compact, fast, and allows effective regularization with the flow priors. We observe that the naive use of cross-camera dense flow priors has a negative effect on the performance, while a careful imposition of reliable cross-camera sparse and within-camera dense flow priors provides a significant boost in performance. Further, we also show that the motion priors are more general and encompass the depth priors in dynamic view synthesis with sparse input views. We demonstrate the effectiveness of our approach on two popular dynamic scene datasets and show that our approach outperforms the state-of-the-art fast and compact dynamic radiance fields by a large margin when only a few viewpoints are available.
@@ Acknowledgements.
This work was supported in part by a grant from Kotak IISc AI/ML centre (KIAC). The first author was supported by the Prime Minister’s Research Fellowship awarded by the Ministry of Education, Government of India.
@@ Appendix A Video Comparisons
We compare our ReFlow-DyRF model with K-Planes (Fridovich-Keil et al., 2023), ablated models, and various other baselines we design. We divide the video comparisons into three sets. In the first set, we compare our model against the vanilla K-Planes and the different baselines we design by incorporating sparse-depth and cross-camera dense flow priors. In the second set, we show the effect of removing each of our priors. Finally, we compare our base DyRF without any priors against K-Planes when dense input views are available. The videos are available on our project website https://nagabhushansn95.github.io/publications/2024/RF-DeRF.html.
@@ Appendix B Performance on Individual Scenes
DyRF 19.82 0.7531 0.3282 0.4619 23.15 0.8361 0.2338 0.3320 25.19 0.8946 0.2018 0.2080 16.93 0.5852 0.4579 0.6320 25.33 0.8819 0.1917 0.3483 25.25 0.8805 0.1909 0.3340 22.61 0.8052 0.2674 0.3860
DyRF + Sparse Depth 19.81 0.7741 0.3088 0.3694 22.64 0.8316 0.2503 0.3332 25.29 0.8813 0.2105 0.1473 16.47 0.5572 0.4732 0.6096 24.59 0.8676 0.2199 0.3325 24.85 0.8783 0.1983 0.2484 22.27 0.7983 0.2768 0.3401
DyRF + Cross-Cam Dense Flow 19.52 0.6990 0.4035 0.6330 23.95 0.8439 0.2443 0.4498 24.94 0.8718 0.2402 0.3664 17.73 0.6043 0.4929 0.7207 25.24 0.8776 0.2055 0.3959 25.24 0.8832 0.1990 0.3937 22.77 0.7966 0.2976 0.4932
ReFlow-DyRF w/o Sparse Flow 20.58 0.8016 0.2923 0.3089 25.61 0.9052 0.1808 0.2936 26.79 0.9174 0.1835 0.2293 17.93 0.6680 0.3944 0.5329 27.07 0.9233 0.1533 0.2365 26.62 0.9099 0.1687 0.2569 24.10 0.8542 0.2288 0.3097
ReFlow-DyRF w/o Dense Flow 21.10 0.8106 0.2863 0.2648 26.03 0.9003 0.1875 0.1791 27.55 0.9225 0.1760 0.1254 19.56 0.7452 0.3622 0.2866 26.85 0.9153 0.1671 0.2060 27.66 0.9280 0.1518 0.1457 24.79 0.8703 0.2218 0.2013
ReFlow-DyRF 21.63 0.8233 0.2700 0.2095 26.29 0.9043 0.1850 0.1601 26.67 0.9135 0.1925 0.1620 20.77 0.7776 0.3194 0.1877 27.95 0.9330 0.1541 0.1320 27.11 0.9192 0.1669 0.1345 25.07 0.8785 0.2146 0.1643
@@
DyRF 19.55 0.8636 0.1533 0.0987 25.93 0.9272 0.1037 0.1344 16.39 0.5377 0.4413 0.2121 16.69 0.5248 0.4144 0.2608 15.02 0.5443 0.3990 0.3938 18.72 0.6795 0.3024 0.2200
DyRF + Sparse Depth 19.25 0.8441 0.1615 0.1149 25.98 0.9249 0.1107 0.2323 15.34 0.5037 0.4566 0.2335 16.02 0.5338 0.4081 0.2832 18.03 0.6691 0.2908 0.2155 18.92 0.6951 0.2855 0.2159
DyRF + Cross-Cam Dense Flow 18.25 0.7215 0.3034 0.2503 24.56 0.8613 0.1802 0.3576 16.20 0.5050 0.4735 0.3198 16.10 0.4545 0.4761 0.2397 17.13 0.6095 0.3677 0.2720 18.45 0.6304 0.3602 0.2879
ReFlow-DyRF w/o Sparse Flow 19.26 0.8339 0.1946 0.1624 25.91 0.9099 0.1245 0.3070 15.73 0.4918 0.4974 0.2781 17.23 0.5346 0.4108 0.2833 17.27 0.6556 0.3138 0.2126 19.08 0.6851 0.3082 0.2487
ReFlow-DyRF w/o Dense Flow 19.51 0.8723 0.1412 0.0666 26.51 0.9397 0.0917 0.0701 17.69 0.5594 0.4154 0.2036 18.45 0.5967 0.3436 0.2112 18.14 0.6901 0.2874 0.0775 20.06 0.7316 0.2559 0.1258
ReFlow-DyRF 19.91 0.8680 0.1449 0.0718 26.55 0.9306 0.1038 0.0953 17.51 0.5749 0.4036 0.1929 18.84 0.6092 0.3471 0.1952 19.24 0.7265 0.2546 0.0639 20.41 0.7418 0.2508 0.1238

### v3

@@ 4.1. Evaluation Setup
Datasets: We evaluate our model on two popular multi-view dynamic scene datasets, namely N3DV (Li et al., 2022) and InterDigital (Sabater et al., 2017) with two, three, and four input views. Following prior work (Fridovich-Keil et al., 2023), we downsample the videos spatially by a factor of two for all the experiments. We use the video at the center of the camera rig for testing and uniformly sample train videos from the remaining videos. N3DV dataset contains six real-world scenes with 17–21 static cameras per scene and 300 frames per viewpoint. The videos have a spatial resolution of $1352\times 1014$ and a frame rate of 30fps. InterDigital dataset contains multiple real-world scenes with 16 static cameras and varying number of frames per scene. We undistort the video frames using the radial distortion parameters provided with the dataset and use the undistorted videos for our experiments. We select five scenes that contain at least 300 frames and choose the first 300 frames. The videos have a spatial resolution of $1024\times 544$ and a frame rate of 30fps spanning 10 seconds in all our experiments.
@@
Table 3. Quantitative results of ablations on N3DV dataset with three input views: We evaluate the significance of the two flow priors we employ, namely the sparse flow prior ($\mathcal{P}_{\text{sf}}$) and the dense flow prior ($\mathcal{P}_{\text{df}}$) by disabling one component at a time. We also show the performance of the DeRF model without any priors for reference.
@@ 4.2. Comparisons and Implementation Details
We mainly compare the performance of our model against K-Planes and HexPlane. We use the official code released by the authors of K-Planes and modify it to implement our model. We refer to our base model without any priors as DeRF model. To test the superiority of our priors, we also compare our model against K-Planes with sparse depth priors and DeRF with dense flow priors across cameras. Since we use two factorized models instead of one used in K-Planes, we reduce the feature dimension in both of our models by half to keep the total number of parameters comparable to K-Planes. We train all the models for 30k iterations on a single NVIDIA RTX 2080 GPU. We randomly pick $s\in\{t-10,t+10\}$ to impose the flow prior losses and set $\lambda_{\text{sf}}=1$ and $\lambda_{\text{df}}=1$. We use the same values as suggested by K-Planes for all the remaining hyperparameters.
@@
Table 4. We test the performance of our DeRF model with dense flow priors across cameras and our reliable flow priors. The performance of the DeRF model (without any priors) is also shown for reference.
@@ 4.3. Results
We show the quantitative performance of our model, K-Planes and HexPlane in Tab. 2, where we observe that our RF-DeRF model outperforms both K-Planes and HexPlane across all the settings on both the datasets. We observe that the performance of all the models is relatively higher on the N3DV dataset as compared to the InterDigital dataset. This is perhaps due to the InterDigital dataset having larger motion and highly textured regions. We also note that the performance of HexPlane is substantially lower than both K-planes and our model. This could be a result of optimizing HexPlane initially with a low-resolution grid, which causes the model to overfit the input-views and not utilize the high-resolution grid in the later stages of training. K-Planes does not suffer from this drawback by using a multi-resolution grid throughout the optimization. From Figs. 6, 7 and 9, we observe that our model is able to correct errors both in moving objects and in static regions over K-Planes. In particular, we observe that K-Planes suffer from blur, deformation, and disappearance of moving objects while rendering novel views, which are mitigated by our model. The improvements are more starkly visible in the supplementary videos. We compare the rendered depth from both the models in Figs. 11 and 11 and observe that our model is able to reconstruct the depth more accurately than K-Planes. Tab. 2 also shows that our DeRF model (without any priors) performs better than K-Planes in the scenes with larger motion, such as the scenes in the InterDigital dataset, perhaps due to the temporal consistency enforced by the canonical volume. This is also visible in the example shown in Fig. 4. In our experiments, RF-DeRF roughly took 1.5 hours and 5GB GPU memory to train on a single scene, and 6 seconds to render a single frame. The size of the saved model parameters is approximately 280MB. Please refer to Fridovich-Keil et al. (2023) for the training time for various models including K-Planes and DyNeRF.
Figure 5. Qualitative examples for depth priors vs flow priors: We observe that sparse depth prior is not very effective in improving the reconstruction quality. However, our sparse flow prior is highly effective in mitigating the distortions, while the use of sparse and dense flow priors in our final model leads to the best reconstruction quality.
Ablations: We analyze the significance of each component of our model by disabling one component at a time in Tab. 3. We observe that the sparse flow prior is the most significant component of our model, giving a large boost in performance when included and a large drop in performance when excluded. The effect of sparse flow prior is more pronounced in terms of depth MAE, which measures the accuracy of the reconstructed 3D scene. This demonstrates the importance of reliable motion priors across cameras, even if sparse. Fig. 8 shows that sparse flow prior is more helpful in mitigating motion related artifacts such as motion blur, while within-camera dense flow prior mainly helps reduce distortions in the static regions. The combination of the two priors leads to the best performance by mitigating both artifacts.
@@
DeRF 0.81 0.39 0.68 0.22
DeRF + $\mathcal{P}_{\text{sd}}$ 0.80 0.34 0.70 0.22
DeRF + $\mathcal{P}_{\text{sf}}$ 0.87 0.20 0.73 0.13
RF-DeRF 0.88 0.16 0.74 0.12
Table 5. We analyze the impact of our sparse flow prior that regularizes both the scene geometry and the learned motion against the sparse depth prior (Deng et al., 2022) that regularizes only the scene geometry. The models are trained with three input views on both the datasets.
Dense cross-camera flow priors vs our priors: To validate our hypothesis that a combination of cross-camera sparse flow and within-camera dense flow priors is more effective than dense flow priors across cameras, we impose cross-camera dense flow priors on our base DeRF model and evaluate the performance on both datasets. From Tab. 4 and Fig. 13, we observe that imposing noisy cross-camera dense flow priors leads to a large drop in performance, whereas our priors help improve the performance significantly.
Depth vs flow priors: We now analyze the effect of flow priors against that of depth priors in dynamic view synthesis with sparse input views. The depth priors guide the model on the exact locations of the objects in the scene, but do not account for the correlation of object locations across time instants. On the other hand, the flow priors account for such inter-frame correlations as well as encompass the information provided by the depth priors in static regions. To experimentally validate our hypothesis, we compare the performance of K-Planes and our DeRF models when regularized with sparse depth priors $\mathcal{P}_{\text{sd}}$ obtained through Colmap sparse reconstruction (Deng et al., 2022). We obtain $\mathcal{P}_{\text{sd}}$ for every frame and use them to constrain the K-Planes model as ${\mathcal{L}_{\text{sd}}}=\|z-\hat{z}\|^{2}$, where $z$ is the depth obtained in Eq. 6 and $\hat{z}$ is the depth prior. From Tab. 5, we observe that sparse flow priors lead to a significant improvement in performance as compared to that of sparse depth priors. The difference in performance can be more clearly observed in Fig. 5 and the supplementary videos. This shows the utility of motion priors in dynamic view synthesis with sparse input views.
Performance with increasing viewpoints: To understand the impact of our priors as the number of input views increases, we compare the performance of RF-DeRF against K-Planes with different number of cameras on a single scene from N3DV dataset in Fig. 13. Our model significantly improves the performance over K-Planes in sparse viewpoint scenarios, while matching the performance of K-Planes with dense input views. Thus, we conclude that our priors are particularly helpful in improving the reconstruction of under-observed regions in the scene.
@@
DeRF 29.95 0.95 0.12 28.14 0.95 0.09
Table 6. We test the performance of our DeRF model against K-Planes with dense input views. Please see the videos in the supplementary for visual comparisons.
Performance with dense input views: We now test the ability of our model in reconstructing the dynamic scene when provided with dense multi-view inputs. Since motion priors may not be needed when dense views are available, we analyze the performance of our base DeRF model (without any priors). From Tab. 6, we observe that our DeRF model is competitive with K-Planes. Thus, one could employ our model in both dense and sparse input scenarios by additionally employing reliable flow priors in the latter case.
@@ 4.4. Limitations and Future Work
Our approach of mapping the scene at every time instant to the canonical volume implies that only the objects present in the canonical volume can be rendered. In other words, our model may not handle objects entering or leaving the scene, which could happen over longer durations. This could be resolved by learning multiple models, each trained on self-contained shorter duration videos. We employ Colmap to generate sparse flow priors in our framework, which is a time-consuming process. In our experiments, we found that the generation of sparse flow priors takes about 45 minutes per scene. There is a need to explore faster approaches to determine reliable sparse correspondences across cameras. Further, our approach requires the cameras to be calibrated and time-synchronized apriori. It would be interesting to optimize the scene, camera parameters, and time-synchronization jointly in a single framework.
@@ 5. Conclusion
We consider the setting of fast dynamic view synthesis when only a few videos of the scene as observed from different static cameras are available. By exploiting the spatio-temporal correlation of the motion field, we design an explicit motion model using factorized representations that is compact, fast, and allows effective regularization with the flow priors. We observe that the naive use of cross-camera dense flow priors has a negative effect on the performance, while a careful imposition of reliable cross-camera sparse and within-camera dense flow priors provides a significant boost in performance. Further, we also show that the motion priors are more general and encompass the depth priors in dynamic view synthesis with sparse input views. We demonstrate the effectiveness of our approach on two popular datasets and show that our approach outperforms the state-of-the-art fast and compact dynamic radiance fields by a large margin when only a few viewpoints are available.
@@ Acknowledgements.
This work was supported in part by a grant from Kotak IISc AI/ML centre (KIAC). The first author was supported partially by the Prime Minister’s Research Fellowship awarded by the Ministry of Education, Government of India.
@@ Appendix A Video Comparisons
We compare our RF-DeRF model with K-Planes (Fridovich-Keil et al., 2023), ablated models, and various other baselines we design. We divide the video comparisons into three sets. In the first set, we compare our model against the vanilla K-Planes and the different baselines we design by incorporating sparse-depth and cross-camera dense flow priors. In the second set, we show the effect of removing each of our priors. Finally, we compare our base DeRF (without any priors) against K-Planes when dense input views are available. The videos are available on our project website https://nagabhushansn95.github.io/publications/2024/RF-DeRF.html.
@@ Appendix B Performance on Individual Scenes
DeRF 19.82 0.7531 0.3282 0.4619 23.15 0.8361 0.2338 0.3320 25.19 0.8946 0.2018 0.2080 16.93 0.5852 0.4579 0.6320 25.33 0.8819 0.1917 0.3483 25.25 0.8805 0.1909 0.3340 22.61 0.8052 0.2674 0.3860
DeRF + Sparse Depth 19.81 0.7741 0.3088 0.3694 22.64 0.8316 0.2503 0.3332 25.29 0.8813 0.2105 0.1473 16.47 0.5572 0.4732 0.6096 24.59 0.8676 0.2199 0.3325 24.85 0.8783 0.1983 0.2484 22.27 0.7983 0.2768 0.3401
DeRF + Cross-Cam Dense Flow 19.52 0.6990 0.4035 0.6330 23.95 0.8439 0.2443 0.4498 24.94 0.8718 0.2402 0.3664 17.73 0.6043 0.4929 0.7207 25.24 0.8776 0.2055 0.3959 25.24 0.8832 0.1990 0.3937 22.77 0.7966 0.2976 0.4932
RF-DeRF w/o Sparse Flow 20.58 0.8016 0.2923 0.3089 25.61 0.9052 0.1808 0.2936 26.79 0.9174 0.1835 0.2293 17.93 0.6680 0.3944 0.5329 27.07 0.9233 0.1533 0.2365 26.62 0.9099 0.1687 0.2569 24.10 0.8542 0.2288 0.3097
RF-DeRF w/o Dense Flow 21.10 0.8106 0.2863 0.2648 26.03 0.9003 0.1875 0.1791 27.55 0.9225 0.1760 0.1254 19.56 0.7452 0.3622 0.2866 26.85 0.9153 0.1671 0.2060 27.66 0.9280 0.1518 0.1457 24.79 0.8703 0.2218 0.2013
RF-DeRF 21.63 0.8233 0.2700 0.2095 26.29 0.9043 0.1850 0.1601 26.67 0.9135 0.1925 0.1620 20.77 0.7776 0.3194 0.1877 27.95 0.9330 0.1541 0.1320 27.11 0.9192 0.1669 0.1345 25.07 0.8785 0.2146 0.1643
@@
DeRF 19.55 0.8636 0.1533 0.0987 25.93 0.9272 0.1037 0.1344 16.39 0.5377 0.4413 0.2121 16.69 0.5248 0.4144 0.2608 15.02 0.5443 0.3990 0.3938 18.72 0.6795 0.3024 0.2200
DeRF + Sparse Depth 19.25 0.8441 0.1615 0.1149 25.98 0.9249 0.1107 0.2323 15.34 0.5037 0.4566 0.2335 16.02 0.5338 0.4081 0.2832 18.03 0.6691 0.2908 0.2155 18.92 0.6951 0.2855 0.2159
DeRF + Cross-Cam Dense Flow 18.25 0.7215 0.3034 0.2503 24.56 0.8613 0.1802 0.3576 16.20 0.5050 0.4735 0.3198 16.10 0.4545 0.4761 0.2397 17.13 0.6095 0.3677 0.2720 18.45 0.6304 0.3602 0.2879
RF-DeRF w/o Sparse Flow 19.26 0.8339 0.1946 0.1624 25.91 0.9099 0.1245 0.3070 15.73 0.4918 0.4974 0.2781 17.23 0.5346 0.4108 0.2833 17.27 0.6556 0.3138 0.2126 19.08 0.6851 0.3082 0.2487
RF-DeRF w/o Dense Flow 19.51 0.8723 0.1412 0.0666 26.51 0.9397 0.0917 0.0701 17.69 0.5594 0.4154 0.2036 18.45 0.5967 0.3436 0.2112 18.14 0.6901 0.2874 0.0775 20.06 0.7316 0.2559 0.1258
RF-DeRF 19.91 0.8680 0.1449 0.0718 26.55 0.9306 0.1038 0.0953 17.51 0.5749 0.4036 0.1929 18.84 0.6092 0.3471 0.1952 19.24 0.7265 0.2546 0.0639 20.41 0.7418 0.2508 0.1238